2. Balance sheet Aa Aa The balance sheet provides a snapshot of the financial condition of a company. Investors and analysts use the information given on the balance sheet and other financial statem ents to make several interpretations regarding the companys financial condition and performance Fuzzy Button Clothing Company is a hypothetical company. Suppose it has the following balance sheet items reported at the end of its first year of operation. For the second year, some parts are still incomplete. Use the information given to complete the balance sheet. Fuzzy Button Clothing Company Balance Sheet for Year Ending December 31 (Millions of Dollars) Year 2 Year 1 Year 2 Year 1 Assets Liabilities and equity Current liabilities: Current assets Cash and equivalents Accounts receivable Inventories $2,767 Accounts payable $0 176 996 $0 0 937 $937 2,813 $3,750 1,266 3,712 $8,437 1,013 Als 2,970 Notes payable Total current assets $6,750 Total current liabilities Net fixed assets: Long-term debt 3,515 $4,687 Net plant and equipment $8,250 Total debt Common equity: 9,141 7,313 3,937 $11,250 $15,000 Common stock Retained earnings Total common equity Total liabilities and equity $14,063 $18,750 Total assets $18,750 $15,000Given the information in the preceding balance sheet-and assuming that Fuzzy Button Clothing Company has 50 million shares of common stock outstanding-read each of the following statements, then identify the selection that best interprets the information conveyed by the balance sheet. Statement #1: Fuzzy Buttons net collection of inventory items increased by more than the firm sold between Years 1 and 2 This statement is because: O Total inventories of raw materials, work-in-process, and final goods decreased by $742 million between Year 1 and Year 2 O The accruals balance decreased by $176 million between Years 1 and 2 O Total inventories of raw materials, work-in-process, and final goods increased from $2,970 milion to $3,712 million between Year 1 and Year 2 Statement #2: Over the past two years, Fuzzy Button Clothing Company has relied more on the use of short-term debt than on long-term debt financing This statement is because: O Fuzzy Buttons total notes payable increased by $59 million, while its common stock account increased O Fuzzy Buttons total current liabilities decreased by $235 million, while its long-term debt account O Fuzzy Buttons total current liabilities increased by $235 million, while its use of long-term debt by $1,828 million decreased by $702 million increased by $702 million Statement #3: One way to interpret the change in Fuzzy Buttons accounts receivable balance from Year 1 to Year 2 is that more customers purchased new items on credit rather than paying off existing credit accounts. This statement is because: O The decrease from $1,266 million to 1,013 million implies a net decrease in accounts receivable and that more customers are paying off their receivables balances than are buying on credit O The 253 increase in accounts receivable means either that Year 1s existing credit customers are not paying off their owed balances and new or existing customers are making additional purchases on credit, or that Year 1s credit customers have repaid their owed balances and Year 2 credit sales have exceeded Year 1s credit sales O The change from $2,970 million to $3,712 million reflects a net accumulation of new credit sales Based on your understanding of the different items reported on the balance sheet and the information they provide, which statement regarding Fuzzy Button Clothing Companys balance sheet is consistent with U.S. Generally Accepted Accounting Principles (GAAP)?
Answer #1

As per rules I am answering the first 4 subparts of the question

1: Net Plant and equipment = Total assets – current assets

= 18750- 8437=10313

2: Cash and cash equivalents = Total current assets-Receivables- Inventories

= 8437-1266-3712 = 3459

3: Total current liabilities = Accounts payable+ Accruals + Notes payable

= 0+ 176+ 996=1172

4: Retained earnings = Total equity – common stock

=14063-9141

=4922